# Grover Speedup from Many Forms of the Zeno Effect

### v4

@@ Abstract
It has previously been established that adiabatic quantum computation, operating based on a continuous Zeno effect due to dynamical phases between eigenstates, is able to realise an optimal Grover-like quantum speedup. In other words, is able to solve an unstructured search problem with the same $\sqrt{N}$ scaling as Grover’s original algorithm. A natural question is whether other manifestations of the Zeno effect can also support an optimal speedup in a physically realistic model (through direct analog application rather than indirectly by supporting a universal gateset). In this paper we show that they can support such a speedup, whether due to measurement, decoherence, or even decay of the excited state into a computationally useless state. Our results also suggest a wide variety of methods to realise speedup which do not rely on Zeno behaviour. We group these algorithms into three families to facilitate a structured understanding of how speedups can be obtained: one based on phase kicks, containing adiabatic computation and continuous-time quantum walks; one based on dephasing and measurement; and finally one based on destruction of the amplitude within the excited state, for which we are not aware of any previous results. These results suggest that there may be exciting opportunities for new paradigms of analog quantum computing based on these effects.
@@ 1 Introduction
It is well known that there are many manifestations of the Zeno effect, for example from measurement, from either dissipative or unitary coupling, or from random phase kicks [1, 2, 3]. Aside from the well-studied case of adiabatic quantum computation [2, 4] which arises effectively from a Zeno effect due to continuous phase kicks, it is less clear how useful these different manifestations will be for directly performing analog quantum computation.
@@
Beyond just being a theoretical construct, demonstrations have been made of the ability of Zeno effects to preserve subspaces, a continuous “partial measurement” on a transmon system [7] was shown to realise a Zeno effect experimentally. Beyond this, quantum Zeno dynamics; dynamics within a non-trivial subsystem isolated using the Zeno effect was demonstrated experimentally within an atomic system [8], with similar effects soon demonstrated for photons [9]. Relatedly, a kind of Zeno effect was demonstrated to be capable of reaching high success probability in an interaction free measurement protocol on transmon systems [10]. For a review of Zeno dynamics, see [11].
@@
If one is to consider analog computing devices based on novel mechanisms, verifying that a mechanism has potential for a speedup is a key step to validate it as a useful tool for quantum computing. In this manuscript, we focus on a speedup on unstructured search. The advantage of this setting, as opposed to general optimisation is that the scaling of the best known possible algorithm is known, so an advantage can be proven. This contrasts with combinatorial optimisation problems, where the best classical scaling is not known. Of particular relevance to this work is the recently proposed entropy computing paradigm of photonic computing [12], which aims to build analog devices making use of a number of effects, including Zeno blockade [13, 14, 15].
Entropy computing is an optical paradigm which can support a fully quantum implementation and has demonstrated promising results in early experiments which make use of classical electronics for feedback [12]. Entropy computing is a paradigm for building photonic based optimisation machines in a way which leverages nonlinear optical effects.
@@
The second key way in which entropy computing differs from the coherent Ising paradigm is the fact that both loss and gain are controlled in the entropy computing paradigm, while only controllable gain is used in the coherent Ising setting. This controllable loss is where the Zeno blockade features in the entropy computing paradigm. By using precisely controlled fabrication techniques, non-linear optics can be used to engineer extremely high photon loss rates, which can lead to Zeno-mediated interactions even at the single photon level. Crucially, in a certain operating regime, a Zeno blockade system can exhibit the underlying effects for one of the computational models we discuss later. Because of the background required, we reserve the details for section 7.
@@
Additionally, Zeno effects have been known to have various applications in quantum computing, for example in this proposal for constructing a universal quantum computer [17]. Therein the authors demonstrate that interaction-free imaging can be extended to the few-atom level, enabling the realization of asymptotically on-demand interaction- and measurement-free quantum logic gates, which are robust against decoherence and detector inefficiency. Interaction-free Zeno gates were experimentally demonstrated in [18] building on methods proposed in [19]. Holonomic quantum gates are often designed and implemented based on adiabatic effects to preserve a subspace [20, 21, 22]. Strong dissipation can also be used to maintain a system within a decoherence free subspace [6]. Furthermore, [5] demonstrated that dissipation based Zeno gates can support universal quantum computation. More recently, [23] showed that engineered dissipative reservoirs can guide many-body quantum systems toward correlated steady states, facilitating scalable entangled many-body states and the exploration of nonequilibrium quantum phenomena. Zeno effects also have applications in error correction [24]. In [25] it was even demonstrated that a non-unitary Zeno effect implemented through measurement can be used to obtain a Grover-type advantage, a specific example of the more general result we show here. Zeno effects have also been found to be useful in implementing constraints in a gate-model setting [26].
@@
More broadly, even beyond Zeno effects, open quantum system effects can have a positive and useful effect on a system. For example, improving aspects such as uniformity in quantum walks [27]. Beneficial effects from open-system interactions have also been directly used in technology, for example the reverse annealing feature as implemented in D-Wave systems only works because of finite temperature dissipation [28, 29]. For another example of the beneficial effect of open system dynamics in that setting, see [30]. Consistent with these trends, our work suggests that there will be many optimal protocols which rely on open system effects but are not directly manifestations of Zeno dynamics.
@@
Additionally, unstructured search is where the exploration between quantum annealing and continuous-time quantum walks began [37]. Continuous time quantum walks now provide an important tool for understanding quantum annealing far from the adiabatic limit. Based on this ability to interpolate in a way that maintained an advantage, there grew the idea that continuous-time quantum walks could be useful for optimisation problems with structure [38] (see also [39] and [40]). From this point it became clear that energetic arguments of the type often used in continuous-time quantum walk could also be applied to multi-stage quantum walks, and this could be taken to a continuum limit which represented rapid quenches [41]. Continuous time quantum walks have since provided a key reference point for understanding diabatic quantum annealing [42, 43, 44, 45]. In this work we hope to provide a similar starting place for theoretical understanding the more general Zeno based hardware being explored in the entropy computing paradigm [12] extensions of the understanding presented here have already begun in [46].
@@ 2.1 Model background
We first define methods which are able to solve the unstructured search problem using a single quantum channel acting on the basis states in the basis defined by the two vectors $\{|\tilde{\omega}\rangle+|m\rangle,|\tilde{\omega}\rangle-|m\rangle\}$ [^1] Normalisation factors have been omitted for brevity in this formula, where $|m\rangle$ is the marked state and $|\tilde{\omega}\rangle$ is a uniform superposition of all other classical basis states defined as
@@
The quantum channel used to solve the problem could for example be applying phases, making a projective measurement, applying decoherence, or allowing the first excited state to dissipate into a computationally useless state we call $|d\rangle$. A crucial step here is to argue over what timescales this quantum channel is physically achievable, as was done in [47] for measurements in this basis. Given that building on the prior work in this paper is crucial to our arguments, we review the key elements of the model from [47] in appendix A. Making arguments about the allowed timescales of these to implement the necessary quantum channels is a crucial step, if not done carefully it could lead to incorrect conclusions. For example, if we assume that dephasing or dissipation can be accomplished in constant time, independent of the size of the system, then this would lead to a system where unstructured search of any size can be accomplished in the same amount of time. Such a result can only come from an unrealistic description of a physical system: any physical quantum system can be efficiently simulated by a universal quantum computer, and it has been proven that such computers cannot exceed the $\sqrt{N}$ scaling of Grover’s algorithm [48], so such a system would have to be “unphysical” or it would contradict a known result. On the other hand, it is not a priori given that every quantum channel should result in any speedup over the classical methods. A major contribution of the current work is to show that for the models we examine, physically realistic versions are able to achieve (but not exceed) $\sqrt{N}$ scaling.
@@
This work builds on a model of unstructured search known as the single avoided crossing model. The model we use was formally introduced in [37], but the underlying concept has been implicit in many other works [33, 49, 34, 50]. This model is relevant to a wide variety of cases of unstructured search, including all computationally relevant qudit models, and for both transverse field (aka hypercube) and complete graph drivers, a full discussion of the relevant search Hamiltonians appears in appendix B.
@@
where $Z=\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right)$, $X=\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right)$, $0\leq\tau\leq 1$ is a dimensionless time parameter, such that $t=\tau\,t_{\mathrm{scale}}g^{-1}_{\mathrm{min}}$ where $t_{\mathrm{scale}}$ defines an overall time scale and $g_{\mathrm{min}}$ is the minimum gap. The function $f(\tau)$ is a control function which defines how the Hamiltonian changes with time with boundary conditions that $f(0)=\infty$ and $f(1)=-\infty$ The prefactor $g_{\mathrm{min}}$ is the minimum spectral gap. This Hamiltonian is mapped such that $|\tilde{\omega}\rangle\rightarrow|1\rangle$ and $|m\rangle\rightarrow|0\rangle$. Within the single avoided crossing model, an optimal annealing schedule is defined as
@@
the results obtained would be identical. Firstly, we can see that the proportion of $\tau$ values where this cutoff is relevant would vanish very quickly it would only be relevant when $\cot^{-1}(g^{-1}_{\mathrm{min}})/\pi>\tau$ or $\cot^{-1}(-g^{-1}_{\mathrm{min}})/\pi<\tau$. Based on the Laurante expansion of the cotangent function, these correspond approximately to $\tau>g_{\mathrm{min}}$ and $\tau<1-g_{\mathrm{min}}$. Since $g_{\mathrm{min}}$ scales as $N^{-\frac{1}{2}}$, the range where this cutoff is relevant quickly vanishes. Furthermore, beyond this cutoff the Eigenstates of the system are to a very good approximation just the marked state and equal superposition. Performing a perturbative expansion, we find that in fact where $f(\tau)\approx g^{-1}_{\mathrm{min}}$, the egenstates of the system will be $\{|m\rangle+O(g^{2}_{\mathrm{min}})|\tilde{\omega}\rangle,|\tilde{\omega}\rangle+O(g^{2}_{\mathrm{min}})|m\rangle\}=\{|m\rangle+O(N^{-1})|\tilde{\omega}\rangle,|\tilde{\omega}\rangle+O(N^{-1})|m\rangle\}$. This suggests further that the error from using $f(\tau)$ instead of the more realistic $\bar{f}(\tau)$ will be of $O(N^{-1})$, which is no larger than the level of error we have already implicitly accepted from ignoring the possibility of randomly guessing the correct solution. For this reason, we can safely proceed with a model based on $f(\tau)$ despite the diverging energy values as we can be assured that the error caused by this approach could at most be as much as neglecting the possibility of random guessing.
@@ 2.2 Invariant Strategy
The single avoided crossing model can generally be invoked for any operation which occurs on a timescale proportional to the inverse of the instantaneous gap $g^{-1}(\tau)$. These certainly include adiabatic and quantum walk protocols, but we will demonstrate that it also includes many other operations. Once time scaling with $g^{-1}(\tau)$ is established, then a family of operations which are invariant with the scaling of $g_{\mathrm{min}}$ can be established, these can either consist of a series of discrete operations, or a continuum quantum channel. Since the purpose of this paper is optimal unstructured search using Zeno dynamics, we focus on the subset of discrete and continuous protocols based on quantum channels which map to Zeno effects (it is however worth noting that our work illuminates many non-Zeno protocols which will also yield an optimal speedup). The final step is to verify that a given protocol yields a success probability which is non-zero,222within the single avoided crossing model, which does not include the exponentially vanishing probability of randomly guessing the correct solution which can be argued either analytically or numerically. The process for constructing a protocol which give an optimal speedup is as follows, operating the protocol with a total runtime $T\propto g^{-1}_{\mathrm{min}}\propto\sqrt{N}$:
@@
    Develop a protocol consisting of a series of implementations of this channel where $H_{\mathrm{ac}}T$ is scale invariant with $g_{\mathrm{min}}$
@@
While all examples here arise from a quantum channel which can achieve $O(1)$ success probability on a timescale proportional to $g^{-1}(\tau)$, it is an interesting hypothetical to consider what other scaling with $g(\tau)$ could mean. Any scaling which is slower than $g^{-1}(\tau)$ would correspond to a speedup better than the known optimal (Grover) speedup, and therefore could only come about from a quantum channel which is somehow unphysical (we give an example later where an unphysical assumption about measurement accuracy appears to show such a speedup). Scaling which is faster than $g^{-1}(\tau)$, but slower than $g^{-2}(\tau)$, would correspond to a non-optimal speedup, since a quantum channel where the runtime scales as $g^{-2}(\tau)$ would result in an overall algorithmic scaling of $N$, which is the best classical scaling. Finally, $g^{-2}(\tau)$ or worse would correspond to no speedup over classical scaling.
We examine the implementation of three types of quantum channels and obtain three families of algorithms, the first family is algorithms where the driving mechanism is phase rotation. This includes the familiar and well understood algorithms of adiabatic quantum computing and continuous time quantum walks. While this family is well understood with already proven speedups we introduce a few minor additions, multi-stage quantum walks and discrete operations we call “phase flipping”.
@@
$\rho\rightarrow|d\rangle\!\langle e|\rho|e\rangle\!\langle d|-\frac{1}{2}|e\rangle\!\langle e|\rho-\frac{1}{2}\rho|e\rangle\!\langle e|.$ (6)

### v5

@@ Abstract
It has previously been established that adiabatic quantum computation, operating based on a continuous Zeno effect due to dynamical phases between eigenstates, is able to realise an optimal Grover-like quantum speedup. In other words, is able to solve an unstructured search problem with the same $\sqrt{N}$ scaling as Grover’s original algorithm. A natural question is whether other manifestations of the Zeno effect can also support an optimal speedup in a physically realistic model (through direct analogue application rather than indirectly by supporting a universal gateset). In this paper we show that they can support such a speedup, whether due to measurement, decoherence, or even decay of the excited state into a computationally useless state. Our results also suggest a wide variety of methods to realise speedup which do not rely on Zeno behaviour. We group these algorithms into three families to facilitate a structured understanding of how speedups can be obtained: one based on phase kicks, containing adiabatic computation and continuous-time quantum walks; one based on dephasing and measurement; and finally one based on destruction of the amplitude within the excited state, for which we are not aware of any previous results. These results suggest that there may be exciting opportunities for new paradigms of analog quantum computing based on these effects.
@@ 1 Introduction
It is well known that there are many manifestations of the Zeno effect for example, from measurement, from either dissipative or unitary coupling, or from random phase kicks [1, 2, 3]. Aside from the well-studied case of adiabatic quantum computation [2, 4] which arises effectively from a Zeno effect due to continuous phase kicks, it is less clear how useful these different manifestations will be for directly performing analog quantum computation.
@@
Beyond just being a theoretical construct, demonstrations have been made of the ability of Zeno effects to preserve subspaces. A continuous “partial measurement” on a transmon system [7] was shown to realise a Zeno effect experimentally. Beyond this, quantum Zeno dynamics; dynamics within a non-trivial subsystem isolated using the Zeno effect was demonstrated experimentally within an atomic system [8], with similar effects soon demonstrated for photons [9]. Relatedly, a kind of Zeno effect was demonstrated to be capable of reaching high success probability in an interaction free measurement protocol on transmon systems [10]. For a review of Zeno dynamics, see [11].
@@
If one is to consider analog computing devices based on novel mechanisms, verifying that a mechanism has potential for a speedup is a key step to validate it as a useful tool for quantum computing. In this manuscript, we focus on a speedup on unstructured search. The advantage of this setting, is that the scaling of the best known possible algorithm is known, so an advantage can be proven. This contrasts with combinatorial optimisation problems, where the best classical scaling is not known. Of particular relevance to this work is the recently proposed entropy computing paradigm of photonic computing [12], which aims to build analog devices making use of a number of effects, including Zeno blockade [13, 14, 15]. Entropy computing is an optical paradigm which can support a fully quantum implementation and has demonstrated promising results in early experiments which make use of classical electronics for feedback [12].
@@
The second key way in which entropy computing differs from the coherent Ising paradigm is the fact that both loss and gain are controlled in the entropy computing paradigm; while only controllable gain is used in the coherent Ising setting. This controllable loss is where the Zeno blockade features in the entropy computing paradigm. By using precisely controlled fabrication techniques, non-linear optics can be used to engineer extremely high photon loss rates, which can lead to Zeno-mediated interactions even at the single photon level. Crucially, in a certain operating regime, a Zeno blockade system can exhibit the underlying effects for one of the computational models we discuss later. Because of the background required, we reserve the details for section 7.
It is worth briefly pointing out that many implementations of coherent Ising machines are based on classical feedback which destroys meaningful quantum superposition [17] in this case one would expect the classical scaling for unstructured search. In fact, simulations can sometimes out-perform the physical machines [18]. Similarly the currently implemented entropy computers use such feedback [12]. Both paradigms however can incorporate quantum search if implemented in an all-optical way [12, 16].
@@
Additionally, Zeno effects have been known to have various applications in quantum computing, for example in this proposal for constructing a universal quantum computer [19]. Therein the authors demonstrate that interaction-free imaging can be extended to the few-atom level, enabling the realization of asymptotically on-demand interaction- and measurement-free quantum logic gates, which are robust against decoherence and detector inefficiency. Interaction-free Zeno gates were experimentally demonstrated in [20] building on methods proposed in [21]. Holonomic quantum gates are often designed and implemented based on adiabatic effects to preserve a subspace [22, 23, 24]. Strong dissipation can also be used to maintain a system within a decoherence free subspace [6]. Furthermore, [5] demonstrated that dissipation based Zeno gates can support universal quantum computation. More recently, [25] showed that engineered dissipative reservoirs can guide many-body quantum systems toward correlated steady states, facilitating scalable entangled many-body states and the exploration of nonequilibrium quantum phenomena. Zeno effects also have applications in error correction [26]. In [27] it was even demonstrated that a non-unitary Zeno effect implemented through measurement can be used to obtain a Grover-type advantage, a specific example of the more general result we show here. Zeno effects have also been found to be useful in implementing constraints in a gate-model setting [28, 29, 30]. Universal bounds for error rates have been derived [31].
@@
More broadly, even beyond Zeno effects, open quantum system effects can have a positive and useful effect on a system. For example, improving aspects such as uniformity in quantum walks [32]. Beneficial effects from open-system interactions have also been directly used in technology. For example, the reverse annealing feature as implemented in D-Wave systems only works because of finite temperature dissipation [33, 34]. For another example of the beneficial effect of open system dynamics in that setting, see [35]. Consistent with these trends. Our work suggests that there will be many optimal protocols which rely on open system effects but are not directly manifestations of Zeno dynamics.
@@
Additionally, unstructured search is where the exploration between quantum annealing and continuous-time quantum walks began [42]. Continuous time quantum walks now provide an important tool for understanding quantum annealing far from the adiabatic limit. Based on this ability to interpolate in a way that maintained an advantage, there grew the idea that continuous-time quantum walks could be useful for optimisation problems with structure [43] (see also [44] and [45]). From this point it became clear that energetic arguments of the type often used in continuous-time quantum walk could also be applied to multi-stage quantum walks, and this could be taken to a continuum limit which represented rapid quenches [46]. Continuous time quantum walks have since provided a key reference point for understanding diabatic quantum annealing [47, 48, 49, 50]. In this work we hope to provide a similar starting place for theoretically understanding the more general Zeno-based hardware being explored in the entropy computing paradigm [12]. Extensions of the understanding presented here have already begun in [51].
@@ 2.1 Model background
We first define methods which are able to solve the unstructured search problem using a single quantum channel acting nontrivially on the basis states in the space defined by the two vectors $\{|\tilde{\omega}\rangle,|m\rangle\}$, where $|m\rangle$ is the marked state and $|\tilde{\omega}\rangle$ is a uniform superposition of all other classical basis states defined as
@@
The quantum channel used to solve the problem could for example be applying phases, making a projective measurement, applying decoherence, or allowing the first excited state to dissipate into a computationally useless state we call $|d\rangle$. A crucial step here is to argue over what timescales this quantum channel is physically achievable, as was done in [52] for measurements in this basis. Given that building on the prior work in this paper is crucial to our arguments, we review the key elements of the model from [52] in appendix A. Making arguments about the allowed timescales of these to implement the necessary quantum channels is a crucial step, if not done carefully it could lead to incorrect conclusions. For example, if we assume that dephasing or dissipation can be accomplished in constant time, independent of the size of the system, then this would lead to a system where unstructured search of any size can be accomplished in the same amount of time. Such a result can only come from an unrealistic description of a physical system: any physical quantum system can be efficiently simulated by a universal quantum computer, and it has been proven that such computers cannot exceed the $\sqrt{N}$ scaling of Grover’s algorithm [53]. Such a system would have to be “unphysical” or it would contradict a known result. On the other hand, it is not a priori given that every quantum channel should result in any speedup over the classical methods. A major contribution of the current work is to show that for the models we examine, physically realistic versions are able to achieve (but not exceed) $\sqrt{N}$ scaling.
@@
This work builds on a model of unstructured search known as the single avoided crossing model. The model we use was formally introduced in [42], but the underlying concept has been implicit in many other works [38, 54, 39, 55]. This model is relevant to a wide variety of cases of unstructured search, including all computationally relevant qudit models, and for both transverse field (aka hypercube) and complete graph drivers. A full discussion of the relevant search Hamiltonians appears in appendix B.
@@
where $Z=\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right)$, $X=\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right)$, $0\leq\tau\leq 1$ is a dimensionless time parameter, such that $t=\tau\,t_{\mathrm{scale}}g^{-1}_{\mathrm{min}}$ where $t_{\mathrm{scale}}g^{-1}_{\mathrm{min}}$ defines an overall time scale and $g_{\mathrm{min}}$ is the minimum gap. The function $f(\tau)$ is a control function which defines how the Hamiltonian changes with time with boundary conditions that $f(0)=\infty$ and $f(1)=-\infty$ The prefactor $g_{\mathrm{min}}$ is the minimum spectral gap. This Hamiltonian is mapped such that $|\tilde{\omega}\rangle\rightarrow|1\rangle$ and $|m\rangle\rightarrow|0\rangle$. Within the single avoided crossing model, an optimal annealing schedule is defined as
@@
the results obtained would be identical. Firstly, we can see that the proportion of $\tau$ values where this cutoff is relevant would vanish very quickly it would only be relevant when $\cot^{-1}(g^{-1}_{\mathrm{min}})/\pi>\tau$ or $\cot^{-1}(-g^{-1}_{\mathrm{min}})/\pi<\tau$. Based on the Laurante expansion of the cotangent function, these correspond approximately to $\tau<g_{\mathrm{min}}$ and $\tau>1-g_{\mathrm{min}}$. Since $g_{\mathrm{min}}$ scales as $N^{-\frac{1}{2}}$, the range where this cutoff is relevant quickly vanishes. Furthermore, beyond this cutoff the eigenstates of the system are to a very good approximation just the marked state and equal superposition. Performing a perturbative expansion, we find that in fact where $f(\tau)\approx g^{-1}_{\mathrm{min}}$, the eigenstates of the system will be $\{|m\rangle+O(g^{2}_{\mathrm{min}})|\tilde{\omega}\rangle,|\tilde{\omega}\rangle+O(g^{2}_{\mathrm{min}})|m\rangle\}=\{|m\rangle+O(N^{-1})|\tilde{\omega}\rangle,|\tilde{\omega}\rangle+O(N^{-1})|m\rangle\}$. This suggests further that the error in the populations from using $f(\tau)$ instead of the more realistic $\bar{f}(\tau)$ will be of $O(N^{-1})$, which is no larger than the level of error we have already implicitly accepted from ignoring the possibility of randomly guessing the correct solution. For this reason, we can safely proceed with a model based on $f(\tau)$ despite the diverging energy values as we can be assured that the error caused by this approach could at most be as much as neglecting the possibility of random guessing. The relative phases will be quite different between the models, but the population is what is relevant for the computation.
@@ 2.2 Invariant Strategy
The single avoided crossing model can generally be invoked for any operation which occurs on a timescale proportional to the inverse of the instantaneous gap $g^{-1}(\tau)$. These certainly include adiabatic and quantum walk protocols, but we will demonstrate that it also includes many other operations. Once time scaling with $g^{-1}(\tau)$ is established, then a family of operations which are invariant with the scaling of $g_{\mathrm{min}}$ can be established, these can either consist of a series of discrete operations, or a continuum quantum channel. Recall that we are interested in the dynamics induced in a two state subspace where the relevant energy scales can be defined by $H_{\mathrm{ac}}$, as discussed in more detail later. Since the purpose of this paper is optimal unstructured search using Zeno dynamics, we focus on the subset of discrete and continuous protocols based on quantum channels which map to Zeno effects (it is however worth noting that our work illuminates many non-Zeno protocols which will also yield an optimal speedup). The final step is to verify that a given protocol yields a success probability which is non-zero,111within the single avoided crossing model, which does not include the exponentially vanishing probability of randomly guessing the correct solution which can be argued either analytically or numerically. Generally any non-trivial dynamics within the relevant subspace should accomplish this goal. The process for constructing a protocol which give an optimal speedup is as follows, operating the protocol with a total runtime $T\propto g^{-1}_{\mathrm{min}}\propto\sqrt{N}$:
@@
    Develop a protocol consisting of a series of implementations of this channel which remains mathematically identical as long as $Tg_{\mathrm{min}}$ is held fixed. In other words $H_{\mathrm{ac}}(\tau)T$ is scale invariant with $g_{\mathrm{min}}$ for all values of unitless time $\tau$.
@@
While all examples here arise from a quantum channel which can achieve $O(1)$ success probability on a timescale proportional to $g^{-1}(\tau)$, it is an interesting hypothetical to consider what other scaling with $g(\tau)$ could mean. Any scaling which is slower than $g^{-1}(\tau)$ would correspond to a speedup better than the known optimal (Grover) speedup, and therefore could only come about from a quantum channel which is somehow unphysical (we give an example later where an unphysical assumption about measurement accuracy appears to show such a speedup). Scaling which is faster than $g^{-1}(\tau)$, but slower than $g^{-2}(\tau)$, would correspond to a non-optimal speedup. A quantum channel where the runtime scales as $g^{-2}(\tau)$ would result in an overall algorithmic scaling of $N$, which is the best classical scaling. Finally, $g^{-2}(\tau)$ or worse would correspond to no speedup over classical scaling.
We examine the implementation of three types of quantum channels and obtain three families of algorithms, the first family is algorithms where the driving mechanism is phase rotation. This includes the familiar and well understood algorithms of adiabatic quantum computing and continuous time quantum walks. While this family is well understood with already proven speedups, we introduce a few minor additions, multi-stage quantum walks and discrete operations we call “phase flipping”.
@@
$\rho\rightarrow|d\rangle\!\langle e|\rho|e\rangle\!\langle d|+|g\rangle\!\langle g|\rho|g\rangle\!\langle g|+|d\rangle\!\langle d|\rho|d\rangle\!\langle d|.$ (6)